OpenStudy (anonymous):

Solve for x 5125 = 5000(1.08)^x

OpenStudy (anonymous):

Check this out: http://www.wolframalpha.com/input/?i=5125+%3D+5000%281.08%29^x

OpenStudy (anonymous):

5125/5000=1.08^x 1.025=1.08^x sqrt(1.025)=1.08. Rest you can solve